Read the directions, show all steps, and answer exactly what should be entered into the box please.
prev Exercise 8.28 Part A Find the energy necessary to put 4 kg initially at rest on Earths surface, into geostationary orbit. Express your answer with the appropriate units. eltte Submit My Answers Give Up
Answer #1

The energy of an object at rest on the earth's surface is

Ui = - GME m/RE

Total mecahnical energy in a circular orbit is

E = U/2 =  - GME m/2r

The energy necessary to put a mass into a cicular geosynchronous orbit with radius is 4.22 * 10^7 m

del E = U/2 - Ui

= - GME m/2r-- GME m/RE

= - GME m(1/RE- 1/2r)

= 6.67 * 10 ^-11( 5.97* 10^24) ( 4 kg) ( 1/6.37 * 10^6 - 1/2( 4.22 * 10^7 m)

=231.17 MJ